# A Survey on Trustworthiness in Foundation Models for Medical Image Analysis

### v1

@@ Abstract
The rapid advancement of foundation models in medical imaging represents a significant leap toward enhancing diagnostic accuracy and personalized treatment. However, the deployment of foundation models in healthcare necessitates a rigorous examination of their trustworthiness, encompassing privacy, robustness, reliability, explainability, and fairness. The current body of survey literature on foundation models in medical imaging reveals considerable gaps, particularly in the area of trustworthiness. Additionally, extant surveys on the trustworthiness of foundation models fail to address their specific variations and applications within the medical imaging domain. This survey paper reviews the current research on foundation models in the major medical imaging applications, with a focus on segmentation, medical report generation, medical question and answering (Q&A), and disease diagnosis, which includes trustworthiness discussion in their manuscripts. We explore the complex challenges of making foundation models for medical image analysis trustworthy, associated with each application, and summarize the current concerns and strategies to enhance trustworthiness. Furthermore, we explore the future promises of these models in revolutionizing patient care. Our analysis underscores the imperative for advancing towards trustworthy AI in medical image analysis, advocating for a balanced approach that fosters innovation while ensuring ethical and equitable healthcare delivery.
Index Terms:
@@
TABLE I: Comparison with existing surveys.
  • $\dagger$ the paper illustrates that different applications have different trustworthiness concerns or solutions
@@
1 Introduction
However, the deployment of foundation models in such a critical sector raises significant concerns regarding their trustworthiness, which encompasses privacy, robustness, reliability, explainability, and fairness (detailed definition see Sec 3). In medical contexts, where decisions have profound implications on patient health, ensuring the trustworthiness of foundation models becomes paramount. It involves rigorous validation against clinical standards, continuous monitoring for performance drift, and mechanisms to interpret model decisions transparently. To fill this gap, we review and discuss recent advances of trustworthiness in foundation models for medical image analysis. To the best of our knowledge, this is the first survey of foundation models for medical image analysis from the trustworthiness perspective, which is different from the existing surveys of foundation models for medical image analysis.
Comparison with existing surveys. We compare existing surveys that discuss similar topics on foundation models, trustworthiness, and particularly those focused on medical imaging. Our survey paper exhibits several distinct advantages over the existing literature, as illustrated in Table I. Existing survey papers on foundation models in medical imaging exhibit significant gaps, particularly in addressing trustworthiness issues. For instance, while Azad et al.  [1], Zhao et al.  [2], and He et al.  [3] discussed foundation models for medical imaging, they fall short in providing detailed analysis on trustworthiness. Whereas, He et al.  [4], Sun et al.  [5], and Liu et al.  [6] survey trustworthiness in foundation models but do not focus on medical imaging applications. Similarly, Salahuddin et al.  [7] and Hasani et al.  [8] focus on trustworthiness issues in medical imaging but do not cover the foundation models. Furthermore, most of the existing survey papers fail to provide a comprehensive examination of trustworthiness and detailed insights into medical imaging application-specific challenges and solutions. In contrast, our survey uniquely integrates an in-depth analysis of trustworthiness across both LLMs and Vision foundation models, emphasizing privacy, robustness, reliability, explainability, and fairness, alongside a comprehensive review of their applications in medical imaging. This holistic approach ensures a more detailed and multifaceted understanding of the current landscape and addresses the critical need for trustworthy AI in medical image analysis, offering valuable insights that are not as extensively covered in other surveys.
Contributions. The contributions of this paper can be highlighted in the following aspects. First, we identify the trustworthiness concerns associated with foundation models for medical image analysis, classifying these issues into five distinct categories: privacy, robustness, reliability, explainability, and fairness. Second, we review the existing literature and categorize them according to the applications of medical image analysis, including segmentation, clinical report generation, medical question and answering (Q&A), and disease diagnosis. Third, we remark on the trustworthiness issues in the existing literature, noting the significant gaps in addressing these concerns. In addition, we propose future research directions, emphasizing the need for innovative approaches to enhance model trustworthiness. This includes advancing privacy techniques, improving model robustness, and fostering fairness and explainability in foundation models.
@@ 1 Introduction
Roadmap. The roadmap of this paper is shown in Figure 1. The subsequent sections are organized as follows: In Section 2, we introduce the background and usage of foundation models for medical image analysis. In Section 3, we define five categories of trustworthiness issues. In Section 4, we categorize the existing literature of foundation models for medical image analysis based on their application, and then by the reported trustworthiness concerns. Finally, we conclude with challenges and future directions in Section 5 before presenting our conclusion in Section 6.
To provide a comprehensive overview, Figure 2 illustrates the landscape of foundation models for medical image analysis studied in this survey. We have covered 31 recent foundation models with applications in medical imaging from 72 research papers, among which 44 are research papers particularly developing for or adapting to medical imaging, published from 2019 to 2024. Given that the type of foundation models is intricately linked to specific medical imaging applications, our survey is systematically categorized according to the prevalent tasks in medical imaging. In the four medical imaging areas we focused on—disease diagnosis, medical Q&A, report generation, and segmentation—each foundation model is associated with multiple medical applications, demonstrating their versatility and applicability across various medical imaging tasks. For instance, models like MedCLIP and ChatGPT are extensively used for disease diagnosis and medical Q&A, while models like SAM [13] and its variants are specialized for Segmentation. LLaMA and Vicuna [14] are used in both medical Q&A and report generation. GPTs [15, 16] are used in all three applications except segmentation. CLIP [17] and its variants are popular choices for medical Q&A and disease diagnosis. In this survey, we also investigate the taxonomy of using foundation models, categorized by training from scratch, fine-tuning, prompt-tuning, or direct off-the-shelf use and how they are linked to trustworthiness aspects. The overlapping and intersecting lines between the use of foundation models and trustworthiness issues indicate the multi-faceted challenges and considerations in deploying these models effectively.
Figure 1: Roadmap of this paper. The subsequent sections of this survey are listed. The icons for section 4 show the input and output data types for each application using foundation models. Tensors represent images.
@@
Figure 2: Landscape of foundation models, their usage, and trustworthiness issues for medical image analysis. The relationship between their usage and trustworthiness issues for different applications is shown. Each thread represents one piece of literature reviewed in this survey.
2 Background and Taxonomy
2-A Definition of Foundation Models
@@
2-B The Use of Foundation Models
@@
2-B1 Training From Scratch
@@ 2-B2 Fine-tuning
2-B3 Prompt-Tuning
Prompt tuning, or prompt engineering, enhances foundation models for specific tasks by crafting input prompts while keeping the model backbone frozen [25]. The terms “prompt tuning” and “prompt engineering” are often used interchangeably, although their focuses are different. Prompt tuning focuses on updating continuous embeddings optimized for specific tasks, while prompt engineering tries out different prompts to find the best ones. Despite these differences, both approaches are commonly employed together in practice to direct the model’s pre-trained knowledge to meet the precise needs of the downstream task, sometimes outperforming full fine-tuning (e.g., VPT [26]). In medical image analysis, well-designed prompts are specific inputs, such as text descriptions that guide the medical foundation model to focus on important features or regions within medical images. For example, a well-designed textual prompt for a medical foundation model analyzing chest X-rays might be, “Focus on the lower lobe of the right lung for potential infiltrates.” This can significantly improve a model’s diagnostic accuracy and provide a resource-efficient way to tailor models for medical diagnostics.
2-B4 Off-the-Shelf
@@
3 Unpacking Trustworthiness in Foundation Models for Medical Image Analysis
3-A Privacy (T1)
Privacy in medical image analysis is critical as patients expect their healthcare service providers to follow necessary safety measures to safeguard their inherent right to the privacy of their information, such as age and sex. This privacy concern can also exist in medical foundation models due to the sensitivity of the health information involved in the extensive training dataset for medical foundation model training [28, 29]. With the increase in data used for training and the complexity of foundation models, there is a growing interest in exploring privacy issues in medical foundation models in both academic and industrial fields.
3-B Robustness (T2)
@@ 3-D Explainability (T4)
3-E Fairness (T5)
Fairness in foundation models refers to the equitable performance of the model across different demographic groups, ensuring that no group experiences significantly lower performance. Similarly, it is also significant to eliminate fairness issues when adapting these foundation models to medical image analysis fields [38]. Due to the common issue of under-representation or imbalance in medical data, fairness in foundation models for medical image analysis is particularly critical. Therefore, it is essential to ensure that these medial foundation models perform consistently well for all populations, mitigating disparate impacts and reducing performance gaps between groups to maintain fairness.
Applications Ref. Foundation Models Usage Trust. Issues Modalities Datasets Body Parts / Organs∗∗
@@ 3-E Fairness (T5)
TABLE II: Papers of foundation models for medical image analysis.
@@
  • X-ray (M1), CT (M2), MRI (M3), PET/CT (M4), Histopathology (M5), Ultrasound (M6), Photography (M7), Endoscopy (M8), Dermoscopy (M9), Microscopy (M10), OCT (M11), Funduscopy (M12)
  • ∗∗ Chest (B1), Head (B2), Neck (B3), Abdomen (B4), Pelvic (B5), Skin (B6), Eye (B7), ENT (B8), Foot (B9), Tooth (B10), Hand (B11), Leg (B12), Knee (B13), Kidney (B14), Porcine Kidney (B15), Spine (B16), Polyp (B17), Brain (B18), Bones (B19), Thorax (B20), Cells (B21), Spleen (B22), Gall Bladder (B23), Esophagus (B24), Liver (B25), Stomach (B26), Aorta (B27), Postcava (B28), Pancreas (B29), Adrenal Gland (B30), Intestine (B31), Bladder (B32), Heart (B33), Prostate (B34), Hip (B35), Wrist (B36), Shoulder (B37)
4 Trustworthiness in Foundation Models for Medical Image Analysis
4-A Overview of the Literature
Table II lists the publications reviewed in this study on the development of medical imaging foundation models, specifically highlighting those that include explicit discussions or evidence regarding their trustworthiness. If several foundation models are used in a paper, the two most representative ones are listed. In addition, the usage of the foundation model is categorized into the four medical imaging applications we focus on, as well as the trustworthiness issues. All the modalities, datasets, and body parts or organs covered in each literature are also identified. Through this structured approach, readers can easily navigate the landscape of current research on medical foundation models and identify key trends and gaps in trustworthiness.
4-B Segmentation
@@
4-B1 Privacy in Segmentation
Privacy concerns in foundation models for medical image segmentation represent a critical challenge, given the sensitive nature of patient data processed by these models. Most foundation models for medical image segmentation are based on SAM architecture [13], which asserts that the 11 M natural image dataset they use is “privacy respecting” - which they specify as meaning that faces and license plates are blurred out. Similarly, in MedSAM [9] the authors leveraged the de-identified public medical images for model training. Federated learning is where model training is done across multiple protected data sources without sharing data [77]. As FedSAM [40] describes, this is a relevant technique for models trained on medical data, as it avoids the patient confidentiality concerns of storing and sharing privacy-sensitive medical data. Despite this, federated learning is rarely used for medical image segmentation models. FedSAM [40] presents a framework for federated SAM fine-tuing across multiple clients, each of which has access to a fraction of the total dataset. Using prostate cancer MRI, brain tumour MRI, nuclei slide images, and fundus photograph sets for training, FedSAM shows comparable performance to SAM, except with the nuclei slide images, where the dataset size is the smallest [40].
4-B2 Robustness in Segmentation
@@
SAMs still lack sufficient robustness on medical images to see widespread use in the medical domain, which has a lower tolerance for uncertainty. A study into the performance of the base SAM on a collection of 556k medical image-mask pairs covering 16 image modalities by Huang et al. noted great performance on some modalities, but partial or complete failure on other image modalities. They state that the performance difference between modalities indicates that SAM “cannot stably and accurately implement zero-shot segmentation on multimodal and multi-object medical datasets” [27]. To address this, Zhang et al. introduces uncertainty estimation for both SAM and MedSAM by using the change in segmentation boundaries as a function of prompt augmentation to generate uncertainty maps [10]. They propose that incorporating uncertainty estimations into SAMs builds trust through better error identification. Jiang et al. notes that features in medical images may have ambiguous boundaries, in contrast to the clear boundaries of most natural images. They propose UA-SAM [45], which integrates a probabilistic model into SAM training and fine-tuning. In addition to fine-tuning the SAM for domains with inherent uncertainty, the adapter in UA-SAM makes the SAM non-deterministic by outputting multiple “plausible” masks for a single input.
@@
Images taken during surgery itself represent an important use case for segmentation. These kinds of images differ in image quality from the natural images used to train segmentation foundation models like SAM. Wang et al. qualitatively assesses the performance of a base SAM on mid-surgery images, and finds that image features like blur or reflections result in inaccurate segmentation [41]. This is concerning as those features would be found throughout the surgery process. Additionally, Wang et al. quantitatively tests the performance of SAM on mid-surgery images from the EndoVis17 image set [78], while imposing 18 different types of noise (brightness, defocus, saturation, JPEG compression etc.) with different severity levels [41]. Some image corruptions, like saturation, only result in a slight decrease in performance, while others, like JPEG compression, cause near total failure [41].
Robustness to different prompt inputs
@@
As mentioned, many SAM-based medical image segmentation models accept two types of prompts: point prompts and box prompts. Prompting gives information to the model on what parts of the image are useful and where boundaries are. This allows a user to contribute to delineating a boundary that the model may not initially detect. This has relevance in the context of medical image segmentation in that the low contrast of medical images (as opposed to the SAM training set of natural images) often requires prompting to get to acceptable performance levels for clinical tasks.
@@
Lei et al. introduces MedLSAM, a foundation model trained on 14,012 CT scan images. MedLSAM can automatically generate box prompts. Evaluating on two sets of CT scan images covering 48 anatomical regions or organs, Lei et al. finds that auto-prompting performance is generally worse than manual prompting, but that for some anatomical features the performance is comparable [46]. Addressing MedLSAM, Zhang et al. proposes automatic prompt generation for SAMs on medical images, a framework they call UR-SAM [10]. UR-SAM perturbs prompts to generate an uncertainty map of potential box prompts. UR-SAM achieves better performance using uncertainty maps than base auto-prompting, though performance is still not as good as manual prompting [10].
The point prompts that SAMs can accept can be positive or negative (containing or not containing the target, respectively). An analysis by Stein et al. gauges the performance benefits of different types of prompts for a base SAM on a set of cardiac magnetic resonance imaging (cMRI) images for segmentation. They find benefits to using positive point prompts, and further benefit in using positive point prompts in conjunction with a box prompt [44]. Interestingly, using positive and negative point prompts in conjunction with a box prompt results in little-to-no improvement over the unprompted SAM. Wang et al. proposes GazeSAM [76], a method for human-computer interaction for SAMs. GazeSAM collects the eye-movement behaviour of users (radiologists), to be used for SAM point prompts. It suggests that prompting with eye-movements can improve the clinician-model relationship by enabling faster prompting [76].
@@
To simulate the errors made by a user during SAM prompting, Roy et al. perform box-prompt segmentation on a collection of CT scan images covering 14 organs while incorporating jitter. The magnitudes of the jitter applied to the box prompts are 0.01, 0.05, 0.1, 0.25, and 0.5. Jitters less than and including 0.1 only marginally decrease performance, whereas jitters of 0.25 and 0.5 cause extremely poor performance. This indicates that some amount of uncertainty in box-prompting can be tolerated by SAM for medical images [43]. Additionally, Roy et al. compares 1 point, 3 point, and 10 points prompts on a multi-organ CT scan image set. Unsurprisingly, the more points that are used, the better the performance. Finally, they compare box prompts to point prompts, and find that box prompts perform considerably better than even the 10 point prompts [43].
@@
Literature in SAM-based medical image segmentation has also explored other types of prompts, including scribble-based [42] and text-based [79], though extensive user testing is not covered or neither across extensive modalities.
4-B3 Fairness in Segmentation

### v2

# Towards Trustworthy Foundation Models for Medical Image Analysis

@@ Abstract
The rapid advancement of foundation models in medical imaging represents a significant leap toward enhancing diagnostic accuracy and personalized treatment. However, the deployment of foundation models in healthcare necessitates a rigorous examination of their trustworthiness, encompassing privacy, robustness, reliability, explainability, and fairness. The current body of survey literature on foundation models in medical imaging reveals considerable gaps, particularly in the area of trustworthiness. Additionally, existing surveys on the trustworthiness of foundation models do not adequately address their specific variations and applications within the medical imaging domain. This survey aims to fill that gap by presenting a novel taxonomy of foundation models used in medical imaging and analyzing the key motivations for ensuring their trustworthiness. We review current research on foundation models in major medical imaging applications, focusing on segmentation, medical report generation, medical question and answering (Q&A), and disease diagnosis. These areas are highlighted because they have seen a relatively mature and substantial number of foundation models compared to other applications. We focus on literature that discusses trustworthiness in medical image analysis manuscripts. We explore the complex challenges of building trustworthy foundation models for each application, summarizing current concerns and strategies for enhancing trustworthiness. Furthermore, we examine the potential of these models to revolutionize patient care. Our analysis underscores the imperative for advancing towards trustworthy AI in medical image analysis, advocating for a balanced approach that fosters innovation while ensuring ethical and equitable healthcare delivery.
keywords:
@@
journal: xxx
[…]
1 Introduction
However, the deployment of foundation models in such a critical sector raises significant concerns regarding their trustworthiness, which encompasses privacy, robustness, reliability, explainability, and fairness (detailed definition see Sec 3). In medical contexts, where decisions have profound implications on patient health, ensuring the trustworthiness of foundation models becomes paramount. It involves rigorous validation against clinical standards, continuous monitoring for performance drift, and mechanisms to interpret model decisions transparently. To fill this gap, we review and discuss recent advances of trustworthiness in foundation models for medical image analysis. To the best of our knowledge, this is the first survey of foundation models for medical image analysis from the trustworthiness perspective, which is different from the existing surveys of foundation models for medical image analysis. The schematic overview in Fig. 1 illustrates how foundation models can be integrated into medical image analysis tasks, while also emphasizing the need to address trustworthiness issues in these applications.
Figure 1: A schematic overview of motivations, foundation model usage, tasks and trustworthiness enhancements discussed in this paper.
Comparison with existing surveys. We compare existing surveys that discuss similar topics on foundation models, trustworthiness, and particularly those focused on medical imaging. Our survey paper exhibits several distinct advantages over the existing literature, as illustrated in Table 1. Existing survey papers on foundation models in medical imaging exhibit significant gaps, particularly in addressing trustworthiness issues. For instance, while Azad et al.  [6], Zhao et al.  [4], and He et al.  [7] discussed foundation models for medical imaging, they fall short in providing detailed analysis on trustworthiness. Whereas, He et al.  [8], Sun et al.  [9], and Liu et al.  [10] survey trustworthiness in foundation models but do not focus on medical imaging applications. Similarly, Salahuddin et al.  [11] and Hasani et al.  [12] focus on trustworthiness issues in medical imaging but do not cover the foundation models. Furthermore, most of the existing survey papers fail to provide a comprehensive examination of trustworthiness and detailed insights into medical imaging application-specific challenges and solutions. In contrast, our survey uniquely integrates an in-depth analysis of trustworthiness across both LLMs and Vision foundation models, emphasizing privacy, robustness, reliability, explainability, and fairness, alongside a comprehensive review of their applications in medical imaging. This holistic approach ensures a more detailed and multifaceted understanding of the current landscape and addresses the critical need for trustworthy AI in medical image analysis, offering valuable insights that are not as extensively covered in other surveys.
@@ 1 Introduction
Table 1: Comparison with existing surveys.
  • 1.
    $\dagger$ the paper illustrates that different applications have different trustworthiness
    concerns or solutions
  • 2.
@@
Contributions. The contributions of this paper can be highlighted in the following aspects.
  1. 1.
    Identifying: We identify trustworthiness concerns in foundation models for medical image analysis, exploring how these issues manifest across different types of foundation model use and various medical imaging tasks.
  2. 2.
    Surveying: We conduct an in-depth review of medical image foundation models in the existing literature and categorize them according to the common applications of medical image analysis that use foundation models, including segmentation, report generation, medical Q&A, and disease diagnosis, and their use of foundation models.
  3. 3.
    Unveiling: We uncover and remark on the trustworthiness issues in the existing literature, highlighting prominent trustworthiness concerns associated with each kind of application, and noting the significant gaps in addressing these concerns.
  4. 4.
    Envisioning: We propose future research directions, emphasizing the need for innovative approaches to enhance model trustworthiness.
Figure 2: Roadmap of this paper.
@@
Roadmap. The roadmap of this paper is shown in Figure 2. The subsequent sections are organized as follows: In Section 2, we introduce the background and usage of foundation models for medical image analysis. In Section 3, we define five categories of trustworthiness issues. In Section 4, we categorize the existing literature of foundation models for medical image analysis based on their application, and then by the reported trustworthiness concerns or strategies for enhancements. Finally, we conclude with challenges and future directions in Section 5 before presenting our conclusion in Section 6.
To provide a comprehensive overview, Figure 3 illustrates the landscape of foundation models for medical image analysis studied in this survey. We have covered 31 recent foundation models with applications in medical imaging from 76 research papers, among which 48 are research papers particularly developing for or adapting to medical imaging, published from 2019 to 2024.
@@
Figure 3: Landscape of foundation models, their usage, and trustworthiness issues for medical image analysis. The relationship between FM usage and trustworthiness issues for different applications is shown. Each thread represents one piece of literature reviewed in this survey.
Given that the type of foundation models is intricately linked to specific medical imaging applications, our survey is systematically categorized according to the prevalent tasks in medical imaging. In the four medical imaging areas we focused on—disease diagnosis, medical Q&A, report generation, and segmentation—each foundation model is associated with multiple medical applications, demonstrating their versatility and applicability across various medical imaging tasks. For instance, models like MedCLIP and ChatGPT are extensively used for disease diagnosis and medical Q&A, while models like SAM [13] and its variants are specialized for Segmentation. LLaMA and Vicuna [14] are used in both medical Q&A and report generation. GPTs [15, 16] are used in all three applications except segmentation. CLIP [17] and its variants are popular choices for medical Q&A and disease diagnosis. In this survey, we also investigate the taxonomy of using foundation models, categorized by training from scratch, fine-tuning, prompt-tuning, or direct off-the-shelf use and how they are linked to trustworthiness aspects. The overlapping and intersecting lines between the use of foundation models and trustworthiness issues indicate the multi-faceted challenges and considerations in deploying these models effectively.
2 Adaptations of Foundation Models in Medical Imaging
2.1 Definition of Foundation Models
@@
2.2 Different Methods for Adapting Foundation Models in Medical Imaging
@@
2.2.1 Training From Scratch
@@ 2.2.2 Fine-Tuning
2.2.3 Prompt-Tuning
Prompt tuning, or prompt engineering, enhances foundation models for specific tasks by crafting input prompts while keeping the model backbone frozen [25]. The terms “prompt tuning” and “prompt engineering” are often used interchangeably, although their focuses differ. Prompt tuning focuses on updating continuous embeddings optimized for specific tasks, while prompt engineering involves trying out different prompts to find the ones that work the best. Despite these differences, both approaches are commonly employed together in practice to direct the model’s pre-trained knowledge to meet the precise needs of the downstream task, sometimes outperforming full fine-tuning (e.g., VPT [26]). In medical image analysis, well-designed prompts are specific inputs, such as text descriptions that guide the medical foundation model to focus on important features or regions within medical images. For example, a well-designed textual prompt for a medical foundation model analyzing chest X-rays might be, “Focus on the lower lobe of the right lung for potential infiltrates.” This can significantly improve a model’s diagnostic accuracy and provide a resource-efficient way to tailor models for medical diagnostics.
2.2.4 Off-the-Shelf
@@
3 Trustworthiness Perspectives in Medical Image Analysis
Trustworthiness in foundation models for medical image analysis is essential due to several critical factors and can cover different perspectives. First, privacy is paramount, as medical images contain highly sensitive patient information requiring robust encryption and secure handling. Secondly, robustness ensures models perform reliably across diverse conditions, minimizing diagnostic errors. Reliability is crucial for consistent, accurate results, avoiding hallucinations and building clinician confidence in the adoption of these tools. Explainability is needed for healthcare professionals to understand and validate the model’s outputs, enhancing trust and safety alignment. Fairness ensures equitable treatment across diverse patient groups, preventing bias and promoting ethical healthcare practices. Although these five perspectives do not cover every trustworthiness concern, they are representative enough to motivate us to survey them. We aim to provide a comprehensive trustworthiness framework that ensures medical foundation models effectively improve patient care and clinical decision-making by addressing these key issues.
3.1 Privacy (T1)
Privacy in medical image analysis is critical as healthcare service providers are expected to follow necessary safety measures to safeguard patients’ private information, such as age and gender. This privacy concern can also exist in medical foundation models due to the presence of health information in the training dataset for medical foundation model training [28, 29]. With the increase in the amount of data used for training and the complexity of foundation models, there is a growing interest in exploring privacy-preserving techniques in medical foundation models in both academia and industry.
3.2 Robustness (T2)
@@ 3.4 Explainability (T4)
3.5 Fairness (T5)
Fairness in foundation models refers to the equitable performance of the model across different demographic groups, ensuring that no group experiences significantly lower performance. Similarly, it is also significant to eliminate fairness issues when adapting these foundation models to medical image analysis fields [38]. Due to the common issue of under-representation or imbalance in medical data, fairness in foundation models for medical image analysis is particularly critical. Therefore, it is essential to ensure that these medical foundation models perform consistently well for all populations, mitigating disparate impacts and reducing performance gaps between groups to maintain fairness.
Drawing from the above various perspectives, trustworthiness in medical image analysis tasks spans a wide range of targets, often dependent on context and application. Additionally, the key concerns about trustworthiness relate to how foundation models are adapted for specific medical imaging applications. There is an urgent need for a systematic review and discussion on the trustworthiness of foundation models in medical image analysis tasks.
Applications Ref. FMs Usage Trust. Modalities Datasets Body Parts / Organs∗∗
@@ 3.5 Fairness (T5)
Applications Ref. FMs Usage Trust. Modalities Datasets Body Parts / Organs∗∗
@@
Table 2: Papers of foundation models for medical image analysis.
  • 1.
@@
  • 3.
    X-ray (M1), CT (M2), MRI (M3), PET/CT (M4), Histopathology (M5), Ultrasound (M6),
    Photography (M7), Endoscopy (M8), Dermoscopy (M9), Microscopy (M10), OCT (M11),
    Funduscopy (M12)
  • 4.
    ∗∗ Chest (B1), Head (B2), Neck (B3), Abdomen (B4), Pelvic (B5), Skin (B6), Eye (B7), ENT (B8),
    Foot (B9), Tooth (B10), Hand (B11), Leg (B12), Knee (B13), Kidney (B14), Porcine Kidney (B15),
    Spine (B16), Polyp (B17), Brain (B18), Bones (B19), Thorax (B20), Cells (B21), Spleen (B22),
    Gall Bladder (B23), Esophagus (B24), Liver (B25), Stomach (B26), Aorta (B27), Postcava (B28),
    Pancreas (B29), Adrenal Gland (B30), Intestine (B31), Bladder (B32), Heart (B33), Prostate (B34),
    Hip (B35), Wrist (B36), Shoulder (B37)
4 Unpacking Trustworthiness in Current Foundation Models Techniques for Medical Imaging Tasks
Table 2 lists the publications reviewed in this study on the development of medical imaging foundation models, specifically highlighting those that include explicit discussions or evidence regarding their trustworthiness. If several foundation models are used in a paper, the two most representative ones are listed. In addition, the usage of the foundation model is categorized into the four medical imaging applications we focus on, as well as the trustworthiness issues. All the modalities, datasets, and body parts or organs covered in each literature are also identified. Through this structured approach, readers can easily navigate the landscape of current research on medical foundation models and identify key trends and gaps in trustworthiness.
4.1 Segmentation
@@
4.1.1 Privacy in Segmentation
Medical image de-identification and federated learning have been adopted in SAM-based models for privacy preservation Privacy concerns in foundation models for medical image segmentation represent a critical challenge, given the sensitive nature of patient data processed by these models. Most foundation models for medical image segmentation are based on SAM architecture [13], which asserts that the 11 M natural image dataset they use is “privacy respecting” - which they specify as meaning that faces and license plates are blurred out. Similarly, in MedSAM [1] the authors leveraged the de-identified public medical images for model training. Additionally, federated learning, a method where model training is done across multiple protected data sources without sharing data [78], have been adopted to train or finetune SAM-based models. It avoids the patient confidentiality concerns of storing and sharing privacy-sensitive medical data Liu et al.  [40] presents a framework, named FedSAM, for federated SAM fine-tuing across multiple clients, each of which has access to a fraction of the total dataset. Using prostate cancer MRI, brain tumour MRI, nuclei slide images, and fundus photograph sets for training, FedSAM shows comparable performance to SAM, except with the nuclei slide images, where the dataset size is the smallest [40]. Similarly, Asokan et al. introduces a method for identifying performance-important layers of a fine-tuned SAM, thus enabling parameter-efficient federated learning, resulting in 6% higher Dice score for a SAM fine-tuned by federated learning on the KiTS19 kidney CT scan dataset [41].
4.1.2 Robustness in Segmentation
@@
Robustness issues across different image modalities. SAM and its variants have claimed to achieve superior performance in image segmentaion, including medical images. Is it a cure-all tool for widespread medical image modalities? Unfortunatley, existing studies show that they still suffer from a lower tolerance for uncertainty. A study into the performance of the base SAM on a collection of 556k medical image-mask pairs covering 16 image modalities by Huang et al. noted great performance on some modalities, but partial or complete failure on other image modalities. They state that the performance difference between modalities indicates that SAM “cannot stably and accurately implement zero-shot segmentation on multimodal and multi-object medical datasets” [27]. To address this, several methods related to uncertainty estimation are proposed to mitigate the robustness issue. Zhang et al. introduces uncertainty estimation for both SAM and MedSAM by using the change in segmentation boundaries as a function of prompt augmentation to generate uncertainty maps [2]. They propose that incorporating uncertainty estimations into SAMs builds trust through better error identification. Jiang et al. notes that features in medical images may have ambiguous boundaries, in contrast to the clear boundaries of most natural images. They propose UA-SAM [46], which integrates a probabilistic model into SAM training and fine-tuning. In addition to fine-tuning the SAM for domains with inherent uncertainty, the adapter in UA-SAM makes the SAM non-deterministic by outputting multiple “plausible” masks for a single input.
@@
Robustness issues on mid-surgery images with blur, reflections, or other types of noise. Images taken during surgery itself represent an important use case for segmentation. These kinds of images differ in image quality from the natural images used to train segmentation foundation models like SAM. Wang et al. qualitatively assesses the performance of a base SAM on mid-surgery images, and finds that image features like blur or reflections result in inaccurate segmentation [42]. This is concerning as those features would be found throughout the surgery process. Additionally, Wang et al. quantitatively tests the performance of SAM on mid-surgery images from the EndoVis17 image set [79], while imposing 18 different types of noise (brightness, defocus, saturation, JPEG compression etc.) with different severity levels [42]. Some image corruptions, like saturation, only result in a slight decrease in performance, while others, like JPEG compression, cause near total failure [42].
@@
Robustness issues on using different prompt inputs. As mentioned, many SAM-based medical image segmentation models accept two types of prompts: point prompts and box prompts. Prompting gives information to the model on what parts of the image are useful and where boundaries are. This allows a user to contribute to delineating a boundary that the model may not initially detect. This has relevance in the context of medical image segmentation in that the low contrast of medical images (as opposed to the SAM training set of natural images) often requires prompting to get to acceptable performance levels for clinical tasks.
@@
Automatic prompting methods are proposed, but the performance is not as good as manual prompting Lei et al. introduces MedLSAM, a foundation model trained on 14,012 CT scan images. MedLSAM can automatically generate box prompts. Evaluating on two sets of CT scan images covering 48 anatomical regions or organs, Lei et al. finds that auto-prompting performance is generally worse than manual prompting, but that for some anatomical features the performance is comparable [47]. Addressing MedLSAM, Zhang et al. proposes automatic prompt generation for SAMs on medical images, a framework they call UR-SAM [2]. UR-SAM perturbs prompts to generate an uncertainty map of potential box prompts. UR-SAM achieves better performance using uncertainty maps than base auto-prompting, though performance is still not as good as manual prompting [2].
@@
The robustness testing on using positive point prompts and box prompts The point prompts that SAMs can accept can be positive or negative (containing or not containing the target, respectively). An analysis by Stein et al. gauges the performance benefits of different types of prompts for a base SAM on a set of cardiac magnetic resonance imaging (cMRI) images for segmentation. Roy et al. perform box-prompt segmentation on a collection of CT scan images covering 14 organs while incorporating jitter. They perform the robustness testing by simulating the errors made by a user, for example, applying different magnitutes jitter to the box prompts. They find that jitters less than and including 0.1 only marginally decrease performance, whereas jitters of 0.25 and 0.5 cause extremely poor performance. This indicates that some amount of uncertainty in box-prompting can be tolerated by SAM for medical images [44]. Additionally, Roy et al. compares using different number of points (i.e., 1 point, 3 point, and 10 points prompts) on a multi-organ CT scan image set. Unsurprisingly, the more points that are used, the better the performance. Finally, they compare box prompts to point prompts, and find that box prompts perform considerably better than even the 10 point prompts [44]. They find benefits to using positive point prompts, and further benefit in using positive point prompts in conjunction with a box prompt [45]. Interestingly, using positive and negative point prompts in conjunction with a box prompt results in little-to-no improvement over the unprompted SAM.
@@
Literature in SAM-based medical image segmentation has also explored other types of prompts, including scribble-based [43], gaze-bsed [77], and text-based [80], though robustness discussion is not covered.
4.1.3 Fairness in Segmentation